# Spectral Mode Enhancement in Coherent-harmonic Dual-comb Spectroscopy Enables Exceeding 300-fold averaging-time reduction

### v3

@@ V Conclusion
In summary, this work has opened up another effective way to enhance the SNR in DCS beyond coherent averaging, through coherent superposition of harmonic pulses. In our proof-of-concept experiments based on temporal phase modulation, we have successfully demonstrated an $\sim$ 18-fold SNR enhancement in a single acquisition, which is equivalent to an SNR gain achieved by coherent averaging over $>$300 periods in conventional DCS. This approach avoids the need for long-time averaging, thus enabling recovery of the inherent excellent temporal resolution of DCS and promising to relax stringent requirements on coherence time and stability of DCS systems. This could potentially have significant implications for future DCS research and applications.

### v4

@@ V Conclusion
In summary, we demonstrate a DCS concept that enhances SNR to reduces coherent averaging time under detector dynamic range limitations by employing phase-pattern higher-repetition-rate frequency combs as sources. We reinterpret the self-imaging process of comb spectra from the perspective of mode interference among sub-pulse trains, and provide a comprehensive description of two architectures for this special form of DCS. In a proof-of-concept experiment, by 20-fold pulse multiplexing, we achieved an $\sim$16.7-fold SNR improvement or a corresponding $\sim$290-fold reduction in averaging time compared to a conventional DCS architecture, in the non-shot-noise-limited regime. These performance gains are expected to decrease to $\sim$4-fold and $\sim$17-fold, respectively, in the shot-noise-limited regime. This approach reduces the reliance on long-term averaging, thereby enabling recovery of the inherent temporal resolution of DCS and potentially relaxing the stringent requirements on coherence time and stability of DCS systems. We expect that this concept can be combined with ultra-high-repetition-rate microcombs for deployment in applications that typically require spectral sampling spacings from hundreds of MHz to the GHz range.